A brass rod is 2m long at a certain temperature.what is its length for a temperature rise of 100kelvin if the expansivity of brass is 18 ×10^-6
1
Expert's answer
2021-06-13T11:29:51-0400

By the definition of the linear thermal expansion, we have:


"\Delta L= L_0\alpha\Delta T,""L-L_0= L_0\alpha\Delta T,""L=L_0(1+\alpha\Delta T),""L=2\ m\cdot(1+18\cdot10^{-6}\ K^{-1}\cdot100\ K)=2.0036\ m."